OpenStudy (anonymous):

Which is an equation of a circle with center (5, 0) that passes through the point (1, 1)? 1.(x - 5)2 + y2 = 17 2.(x + 5)2 + y2 = 17 3.(x - 1)2 + (y - 1)2 = 17 4.(x + 5)2 + y2 = 24

OpenStudy (anonymous):

1.

OpenStudy (scorcher219396):

The general graphing formula of a circle is (x-h)^2+(y-k)^2=r^2 where (h,k) is the center and r is the radius So if it tells you that the center is at (5,0), then you know that the left side of the equation is (x-5)^2+y^2 For the right, you need to know what the radius is If it passes through (1,1), then you know that the radius is the distance between (1,1) and (5,0) If you set up a triangle between those points, the long side would be 4, the short side 1, and using the pythagorean theorem you would get sqrt(17) as the radius Now you can put that in the formula as well, and have (x-5)^2+y^2=17, which is answer 1